# Proprioceptive State Estimation for Amphibious Tactile Sensing

### v1

@@ 5.3 Deformation Estimation using Touch Haptic Device
Figure 5: Estimated deformation field of the soft finger using the proprioceptive shape estimation method. (A) The Touch haptic device is used to make contact with the soft finger at different locations, while at the same time, the ground-truth positions and the reconstructed positions of contact points are recorded. (B) The pushing path points for three-dimensional and corresponding error of three distinct contact locations during the experiment. (C) The pen-nib of the touch haptic device is commanded to be pushed forward five times at fifty locations on the half sides of the soft finger. The map of the mean error norm is interpolated using fifty sampled contact locations. (D) The distribution of corresponding total error norm along the Z axis of the soft finger. (E) The corresponding total error norm of sampled contact points.
As is shown in Fig. 5A, the pen-nib of a Touch haptic device is used to stick tightly to some contact point on the soft finger and push forward five times during the experiment. We randomly sample fifty contact points covering half of the soft finger’s sides and record the pen-nib’s position and the corresponding point on the estimated deformed mesh model.
@@
Similar to the calibration process of the experiment using a motion capture system, we solve the barycentric coordinates in Eq. 16, using the initial sticking position of pen-nib and the undeformed vertex position of the tetrahedron nearest to the sticking point. Since there is no slipping between the contact point and the pen-nib, recording the pushing position of the pen-nib for a randomly selected point is equivalent to collecting the ground truth deformation field of the soft finger evaluated at that point. Three pushing path points and the corresponding error between estimation and ground truth for three randomly selected contact locations are demonstrated in Fig. 5B. The pushing duration lasts ten seconds for each selected contact location, and the data is recorded at a sampling frequency of 20 Hz. Due to the variation among the pushing path for three contact locations, the error distribution exhibits slightly different, but all lie in the $(0,2.5)$ mm range.
@@
Compared with the experiment using six motion capture markers for shape estimation evaluation, the measurement points using the Touch device are dense and have more extensive coverage of the soft finger, revealing the spatial error distribution of the estimation method. As is shown in Fig. 5C, mean errors of deformation estimation evaluated at fifty randomly selected contact locations are visualized. Two side views of continuous error distribution for the soft finger are interpolated with errors of fifty sampled locations using a Gaussian-kernel-based nearest neighbor method [63].
[…]
5.4 Underwater Exteroceptive Estimation of Object Shape
@@
Figure 6: Underwater shape estimation of a vase using proprioceptive state estimation of the soft finger. (A) Experimental setup for underwater shape estimation, where a Robotiq Hand-E gripper, installed with two proprioceptive soft fingers and an extension link, is mounted on a Franka Emika Panda robot arm. The gripper is programmed to perform a series of actions periodically, including gripping, releasing, and moving along the x-axis for a fixed distance. At the same time, a vase is fixed at the bottom of the tank in the lab. (B) Contact surface patch prediction using Gaussian process implicit surface (GPIS) with the soft finger. (C) Experiment pipeline for underwater shape estimation of a vase. (D) Evaluation of the reconstructed vase shape on some cutting sectional planes, measured in Chamfer Distance.
@@
As is shown in Fig. 6A, we mounted the integrated soft finger on the Robotiq’s Hand-E gripper and attached the whole design at the end of a Franka Emika Panda robotic manipulator with an extension link. The gripper has an ingress protection rating of IP67, and we verified its consistent performance for temporary submergence under the water between 0.15 m and 1 m for 30 minutes. Our previous work extensively tested the gripper’s underwater servoing capabilities for reactive grasping [64]. In this study, we use the same gripper for underwater object shape estimation in a lab tank. For more intensive usage underwater in the field, one can always replace the Hand-E gripper with a more appropriate one.
With the gripper submerged underwater, the grasping system is programmed to sequentially execute a series of actions, including gripping the object, releasing the object, and moving along a prescribed direction for a fixed distance for underwater object shape information acquisition. See Movie S3 in the Supplementary Materials for a video demonstration.
The experiment setup is shown in Fig. 6B(i). By mounting the target object at the bottom of the tank, we assume:
  • The object’s pose is fixed and calibrated with the gripper.
  • Passive object shape exploration is considered for object coverage.
@@
The inference of the GPIS model is computationally intractable for the large $N$ measurement that accrues from high-dimensional tactile measurements. Instead of predicting the whole object surface by accumulating all the collected data, we only query a local GPIS model approximated using current observed contact data in a local focus area and build the surface incrementally.
@@
Local GPIS Model Inference: A training set containing contact interface points $\mathbf{x}_{c}$ and corresponding augmented control points are collected each time a grasping action is performed. Before querying the local GPIS model in the interested area, hyper-parameters $\sigma^{2}_{f}$ and $l$ associated to Eq. 13 are optimized first using the standard training method for Gaussian processes, i.e., maximizing the marginal likelihood. Then, we evaluate the local GP on voxel grid points at a resolution of 0.2 mm in the interested area and keep those points with zero mean of Eq. 14 as estimated points on the surface patch of the object.
@@
Global Registration of Local Patches: With the calibrated object pose w.r.t. the gripper, we programmed the grasping system to follow a pre-defined path for object shape exploration. As is shown in Fig. 6B(iv), each time after GPIS query in the local 3D region, a global registration action is performed by transforming these local iso-surface points into the global space. Leveraging the continuous nature of the pre-defined exploration path, a simple surface registration strategy is used, where only the points of the estimated surface patch corresponding to moving distance are kept, and points of overlapping intervals belonging to the latest estimated surface patch are rejected. As is shown in Fig. 6C, after initialization of the relative pose between the gripper and the object, the shape of the object is continuously reconstructed using the described passive exploration strategy.
@@
Object Shape Estimation Evaluation: In Fig. 6D, we present our method on real data collected during the underwater tactile exploration experiment. The shape estimates at each cutting sectional plane are compared concerning the ground truth using the Chamfer Distance (CD) [65], a commonly-used shape similarity metric. We chose five vertical cutting planes and one horizontal sectional plane for reconstructed object surface evaluation. For each cutting plane, a calibration error exists between the vase and the Hand-E gripper, leading to the expected gap between the reconstructed and ground truth points. In addition to the systematic error, we have observed a slight decrease in the CD metric values between planes 1 and 5 compared to planes 2, 3, and 4, which could be attributed to the limitations of the soft finger in adapting to small objects with significant curvature. On the other hand, by employing tactile exploration actions with a relatively large contact area on the soft finger’s surface, the shape estimation of objects similar in size to the vase can be accomplished more efficiently, typically within 8-12 touches. The 3D-printed vase has dimensions of approximately 80 mm by 80 mm by 140 mm.
@@ 6 Discussion
6.1 Sensorization of Soft Tactile Finger
Soft robots hold immense potential for delicate and controlled manipulation tasks, aiming to mimic the dexterity and adaptability of human hands [66]. Integrating high-resolution tactile sensing into such systems dramatically enhances their ability to perceive and understand the physical properties of objects they handle. However, equipping soft robots with comparable proprioceptive capabilities poses significant challenges, particularly for robotic fingers, which constantly interact with the environment and operate in confined spaces [67].
To overcome these challenges, researchers have explored various approaches. One strategy involves implementing soft and stretchable sensors that conform to the finger’s shape [68], while another embeds sensors directly within the finger’s structure. Vision-based tactile sensing presents a promising alternative, offering access to high-resolution contact information through an embedded camera [69]. Recent studies have successfully demonstrated the feasibility of this approach on rigid finger models [70], prompting exploration of its integration with compliant soft fingers [71, 72].
This study builds upon our previous work on soft finger design and in-finger camera integration to achieve underwater proprioceptive shape sensing. We developed a robust and cost-effective soft tactile finger design with transferable capabilities for underwater robotics applications by employing a simple waterproofing technique for the camera board.
6.2 Morphological Perception with Soft Robots
Morphological perception using soft bodies is a concept that involves utilizing the deformable nature of soft materials to enhance perception capabilities in robots [73]. Traditional rigid robots often struggle to adapt to complex environments and gather rich sensory information. Soft robots, on the other hand, leverage the deformability of soft materials to enhance their perception capabilities. These robots can actively change their shape and morphology by employing compliant materials or structures, interacting with the environment more nuancedly. This ability to deform in response to external stimuli becomes a powerful tool for perception [74]. It allows robots to “feel” their surroundings passively and gather information about object shape, texture, and internal composition [75].
Researchers have explored various soft robotic morphologies for object shape estimation. One approach utilizes curve-like tactile whiskers that bend and strain upon contact, providing information about the object’s contour [76]. Another strategy employs flat elastomer pads that conform to the object’s surface, estimating its shape through the pad’s deformation [77]. Dome-shaped fingertips represent another morphology, using contact area and pressure distribution to infer shape and texture [78]. Each of these morphologies offers unique advantages and capabilities for accurate and versatile object shape estimation.
In this study, we propose a method that leverages the known deformation model of a specific soft finger to reconstruct its changing morphology during contact. The robot can make accurate proprioception state estimations using relatively sparse observations by understanding how the finger deforms based on its internal structure and material properties. This enhanced ability to perceive and interact with objects opens up exciting possibilities for robots operating in diverse and unpredictable environments.
7 Conclusion, Limitations, and Future Works
This paper presents a novel method for estimating a soft finger’s deformation and contact surface using visual observations and a model-based approach. This “constrained geometry optimization” technique achieves efficient deformation reconstruction, comparable to traditional methods (e.g., ABAQUS) and modern alternatives (e.g., ARAP). Our method shows impressive speed, running nearly 40 times faster than ABAQUS while maintaining similar accuracy. Experimental validation demonstrates a median error of just 1.96 mm (2.1% of finger length) in overall body deformation estimation. Beyond deformation, we detail a promising pipeline for estimating the surface of contacted objects using information about the finger’s shape under pressure. This application is showcased underwater with soft tactile fingers integrated into a waterproof gripper.

### v2

@@ 4.1.3 Deformation Estimation using Touch Haptic Device
Figure 5: Estimated deformation field of the soft finger using the proprioceptive state estimation method. (A) The Touch haptic device is used to make contact with the soft finger at different locations while simultaneously recording the ground-truth positions and the reconstructed positions of contact points. (B) Three sampled pushing trajectories of the pen-nib and corresponding measurements from the proprioceptive state estimation method. Total Errors are reported in the last column. The pen-nib of the touch haptic device is pushed forward and backward five times at each location. (C) The fifty testing locations sampled are spread over half of the side of the soft finger. The mean error norm map is interpolated using the values of the fifty sampled contact locations. (D) The distribution of the total errors along the height (Z-axis) of the soft finger. (E) The distribution of the total errors of sampled contact points.
@@
Similar to the calibration process when using the motion capture system, we solve the barycentric coordinates in Eq. (16) using the initial contact position of pen-nib and the undeformed vertex position of the tetrahedron nearest to the contact point. Since there is no slipping between the contact point and the pen-nib, recording the pushing position of the pen-nib for a randomly selected point is equivalent to collecting the ground truth deformation field of the soft finger evaluated at that point. Figure 5B shows three selected pushing trajectories and the corresponding errors between estimation and ground truth. The pushing duration lasts around ten seconds for each location and is rescaled to 1 in the plot. The data is recorded at 20 Hz. Due to the variations among the pushing trajectories among the three locations, the errors are slightly different, but all lie within a 2.5 mm range.
@@
The haptic device measurements cover an extensive portion of the soft robotic finger, revealing further details regarding the spatial distribution of the estimation errors. We visualize the mean errors of deformation estimation evaluated at the fifty randomly selected contact locations in Fig. 5C. We interpolated two side views of continuous error distribution for the soft robotic finger with errors of all sampled locations using a Gaussian-kernel-based nearest-neighbor method [70].
Contact locations near the observed AMH constraint are expected to exhibit fewer errors due to penalized computation near this region during deformation optimization. We plot the error distribution of all sampled locations along the $Z$ axis in Fig. 5D. Contact locations with a similar height to the AMH constraint exhibit a smaller and more concentrated error distribution. Figure 5E shows the error histogram of the overall experiment records, where the median of estimated error for the whole-body deformation is 1.96 mm, corresponding to 2.1% of the finger’s length. (See Movie S2 in the Supplementary Materials for a video demonstration.)
4.2 On Amphibious Tactile Sensing for PropSE
Here, we further investigate our proposed method in amphibious tactile sensing through three experiments in lab conditions. We begin by benchmarking our proposed VBTS method at controlled turbidity underwater. Then, we present a touch-based object shape reconstruction task to demonstrate the application of our proposed solution for amphibious tactile sensing. Finally, we present a full-system demonstration by attaching our robotic finger to the gripper of an underwater Remotely Operated Vehicle (ROV) for underwater grasping in a water tank, which we plan to implement further in the field test soon.
4.2.1 Benchmarking VBTS Underwater against Turbidity
Our proposed rigidity-aware AMH method effectively transforms the visual perception process for deformable shape reconstruction into a marker-based pose recognition problem. Therefore, the benchmarking of our vision-based tactile sensing solution underwater is directly determined by successfully recognizing the fiducial marker poses used in our system under different turbidity conditions. Turbidity is an optical characteristic that measures the clarity of a water body and is reported in Nephelometric Turbidity Units (NTU) [71]. It influences the visibility of optical cameras for underwater inspection, inducing light attenuation effects caused by the suspended particles [72]. As one of the critical indicators for characterizing water quality, there have been rich studies on the turbidity of large water bodies worldwide. For example, previous research [73] shows that the Yangtze River’s turbidity is measured between 1.71 and 154 NTU.
Figure 6: Benchmarking results in different turbidity conditions underwater in a lab tank. (A) The experiment was set up in a room with controlled ambient lighting of 3,000 lumens placed atop the tank (not shown in this picture). (B) Images taken by adding condensed standard turbidity liquid to increase the water turbidity from 0 to 160 NTU, including i) experiment pictures taken by an external camera at the same angle as (A); ii) raw images captured by the in-finger vision overlayed with triad coordinates to indicate successful pose recognition; and iii) digitally enhanced images overlayed with triad coordinates to indicate successful pose recognition. (C) Results on the pose recognition success rate of the ArUco marker from the in-finger vision under increasing tank turbidity when pushing the soft robotic finger at different target displacements, with or without image enhancement.
We investigated the robustness of our proposed VBTS solution in different water clarity conditions by mixing condensed standard turbidity liquid with clear water to reach different turbidity ratings. Figure 6A shows the experiment setup. Our proposed soft robotic finger is installed on a linear actuator in a tank filled with 56 liters of clear water. A probe is fixed under the soft robotic finger, inducing contact-based whole-body deformation when the finger is commanded to move downward. The tank is placed in a room with controlled ambient lighting of 3,000 lumens placed atop the tank. We controlled the linear actuator for each turbidity condition so that the finger moved downward along the $x$ axis. This enabled us to record the ArUco image streams when fixed 0, 2, 4, 6, and 8 mm displacements in $D_{x}$ are reached. For example, the three images shown in the first column of Fig. 6B are i) the experiment scenario taken at the same angle as Fig. 6A when the turbidity is zero (before adding condensed standard turbidity liquid), ii) a sample of the raw image captured by our soft robotic finger’s in-finger camera, and iii) image enhancement based on the image shown in ii), respectively. The water tank’s clarity is modified by adding specific portions of condensed standard turbidity liquid to reach different turbidity ratings at 10 NTU per step (images for 20 NTU per step increase are shown in Fig. 6B for the ease of visualization), increasing from 0 to 160 NTU, covering the Yangtze River’s turbidity range.
For each of the $D_{x}$ positions, we recorded 1,000 images using our soft robotic finger’s in-finger camera to obtain the pose recognition success rate (%) under each turbidity rating, before or after image enhancement, reported in Fig. 6C. The results reported in Fig. 6C involve data aggregated from 85,000 raw images (1,000 images per NTU step per ArUco position $\times$ 17 NTU steps $\times$ 5 ArUco positions) from in-finger vision for ArUco pose recognition, which is doubled after image enhancement, resulting a total of 170K images.
In our experiment, for the turbidity range between 0 and 40 NTU, the raw images captured by our in-finger vision achieved a 100% success rate in ArUco pose recognition. At 50 NTU turbidity, the first sign of failed marker pose recognition was observed when the most considerable deformation was induced at $8$ mm of $D_{x}$. Our experiment shows that this issue can be alleviated using simple image enhancement techniques to regain a 100% marker pose recognition success rate. However, the marker pose recognition performance under large-scale whole-body deformation quickly deteriorated when the turbidity reached 60 NTU and eventually became unusable at 70 NTU. Image enhancement could effectively increase the upper bound to 100 NTU to reach an utterly unusable marker pose recognition in large-scale whole-body deformation. However, for small or medium whole-body deformations measured by $D_{x}\leq 6$ mm, our system remains functional until around 100 NTU in turbidity, where simple image enhancement techniques help for a balanced consideration between algorithmic cost, engineering complexity, and system performance.
@@ 4.2.1 Benchmarking VBTS Underwater against Turbidity
Figure 7: Underwater shape estimation of a vase using proprioceptive state estimation of the soft finger. (A) This is the experimental setup for underwater shape estimation. A Robotiq Hand-E gripper, installed with two proprioceptive soft fingers and an extension link, is mounted on a Franka Emika Panda robot arm. The gripper is programmed to perform a series of actions periodically, including gripping, releasing, and moving along the x-axis for a fixed distance. At the same time, a vase is fixed at the bottom of the tank in the lab. (B) Contact surface patch prediction using Gaussian process implicit surface (GPIS) with the soft finger. (C) Experiment pipeline for underwater shape estimation of a vase. (D) Evaluation of the reconstructed vase shape on some cutting sectional planes, measured in Chamfer Distance.
For turbidity above 100 NTU, simple image enhancement provides limited contributions to our system. Our experiment shows that when the turbidity reached 160 NTU, our in-finger system failed to recognize any ArUco pose underwater, even after image enhancement. Since blurry images of the marker remain visible in the captured images, we can 1) use more advanced image processing algorithms, 2) use better imaging hardware, 3) apply stronger ambient lighting, and 4) redesign the marker pattern specifically for underwater usage to systematically increase the upper bound of the turbidity rating for marker-based posed estimation in contact-based amphibious grasping using vision-based tactile sensing methods. Results obtained from this experiment provide a general understanding of the potential regions for amphibious grasping characterized by turbidity with possible solutions to improve further.
4.2.2 Underwater Exteroceptive Estimation of Object Shape
@@
In this experiment, we apply our soft robotic finger with in-finger vision to a contact-based shape reconstruction task to demonstrate our solution’s capabilities in underwater exteroceptive estimation. Shown in Fig. 7A is the experimental setup conducted in the lab condition using the same water tank as the one used in the previous experiment. In this case, we used a parallel two-finger gripper (HandE from Robotiq, Inc.) attached to the wrist flange of a robotic manipulator (Franka Emika) through a 3D-printed cylindrical rod for an extended range of motion. Our soft robotic fingers are attached to each fingertip of the gripper through a customized adapter fabricated by 3D printing. Our previous work extensively tested this IP67 gripper’s underwater servoing capabilities for reactive grasping during temporary submergence under the water [74]. In this study, we use the same gripper for underwater object shape estimation in a lab tank. One can always replace the Hand-E gripper with a professional underwater gripper for more intensive underwater usage in the field.
@@
With the gripper submerged underwater, the system is programmed to sequentially execute a series of actions, including gripping and releasing the object and moving along a prescribed direction for a fixed distance to acquire underwater object shape information, as shown in Fig. 7B(i). By mounting the target object at the bottom of the tank, we assume that 1) the object’s pose is fixed and calibrated with the gripper and 2) passive object shape exploration is considered for object coverage. The inference of the GPIS model is computationally intractable for the large $N$ measurement that accrues from high-dimensional tactile measurements. Instead of predicting the whole object surface by accumulating all the collected data, we only query a local GPIS model approximated using current observed contact data in a local focus area and build the surface incrementally, as shown in Figs. 7B(ii)&(iii).
Local GPIS Model Inference
@@
A training set containing contact interface points $\mathbf{x}_{c}$ and corresponding augmented control points are collected each time a grasping action is performed. Before querying the local GPIS model in the interested area, hyper-parameters $\sigma^{2}_{f}$ and $l$ associated to Eq. (13) are optimized first using the standard training method for Gaussian processes, i.e., maximizing the marginal likelihood. Then, we evaluate the local GP on voxel grid points at a resolution of 0.2 mm in the interested area and keep those points with zero mean of Eq. (14) as estimated points on the surface patch of the object.
Local Patches Concatenation
@@
After calibrating the object pose to the gripper, we programmed the grasping system to follow a pre-defined path for object shape exploration. As is shown in Fig. 7B(iv), each time after GPIS query in the local 3D region, a global registration action is performed by transforming these local iso-surface points into the global space. Leveraging the continuous nature of the pre-defined exploration path, a simple surface concatenation strategy is used, where only the points of the estimated surface patch corresponding to moving distance are kept, and points of overlapping intervals belonging to the latest estimated surface patch are rejected. As is shown in Fig. 7C, after initialization of the relative pose between the gripper and the object, the shape of the object is continuously reconstructed using the described passive exploration strategy.
Object Shape Estimation Evaluation
@@
In Fig. 7D, we present our method on actual data collected during the underwater tactile exploration experiment. The shape estimates at each cutting sectional plane are compared concerning the ground truth using the Chamfer Distance (CD) [75], a commonly-used shape similarity metric. We chose five vertical cutting planes and one horizontal sectional plane for reconstructed object surface evaluation. For each cutting plane, a calibration error exists between the vase and the Hand-E gripper, leading to the expected gap between the reconstructed and ground truth points. In addition to the systematic error, we have observed a slight decrease in the CD metric values between planes 1 and 5 compared to planes 2, 3, and 4, which could be attributed to the limitations of the soft finger in adapting to small objects with significant curvature. On the other hand, by employing tactile exploration actions with a relatively large contact area on the soft finger’s surface, the shape estimation of objects similar in size to the vase can be accomplished more efficiently, typically within 8-12 touches. The 3D-printed vase has dimensions of approximately 80 mm by 80 mm by 140 mm. (See Movie S3 in the Supplementary Materials for a video demonstration.)
4.2.3 Vision-based Tactile Grasping with an Underwater ROV
Here, we provide a full-system demonstration by using our vision-based soft robotic fingers on a underwater Remotely Operated Vehicle (ROV, FIFISH PRO V6 PLUS by QYSEA [^2] https://www.qysea.com/). It includes a single-DOF robotic gripper, which can be modified using the proposed soft fingers with customized adaptors.
The experiment results reported in Section 4.2.1 already benchmark our system’s promising capabilities for real-time underwater tactile sensing. As shown in Fig. 6B, the water at 20 NTU or above is already very challenging to clearly observe from a third-personal perspective. Experiments underwater would require additional cost to prepare a second underwater ROV to record videos when the water is clear enough. However, as analyzed above, our in-finger vision could perform nicely at a much higher NTU range. Therefore, in this section, we only conducted this experiment in a lab tank to demonstrate our system’s integration with an existing underwater ROV system during an underwater task.
Figure 8: Demonstration of our soft robotic finger with in-finger vision for tactile sensing underwater. (A) Key components involved in the test. (B) Screenshot of a 4K image captured by the underwater ROV’s onboard camera when our fingers are holding a conch after successfully grasping. (C) & (D) A screenshot of the images captured by an in-finger vision camera in the left and right fingers while holding the conch. (E) & (F) Whole-body deformation reconstruction for both fingers based on the images captured by the in-finger vision cameras, respectively.
Shown in Fig. 8A is a brief overview of the system and the scene. Our fingers are attached to the underwater ROV’s gripper through 3D-printed adaptors to replace the default rigid fingers. Our design conveniently introduced omni-directional adaptation capability to the gripper’s existing functionality with added capabilities in real-time tactile sensing underwater. Shown in Fig. 8B is a screenshot of the image taken by the ROV’s onboard camera, recording 4K videos in real-time. In this experiment, both soft robotic fingers are installed with in-finger vision, capturing images shown in Figs. 8C&D. Using these in-finger images, we can use the methods proposed in this work to achieve real-time reconstruction of contact events on our soft robotic finger in Figs. 8E&F, while performing grasping tasks underwater.
See Movie S4 in the Supplementary Materials for a video demonstration. Besides the capabilities demonstrated in this paper, we also identified an interesting observation during the experiment, adding to the benefits of having soft robotic fingers for underwater ROVs compared to the traditional rigid ones. While performing grasping underwater, the target objects are usually at the bottom. It is challenging for the underwater ROV to approach the target object smoothly and slowly, even in the lab tank with no water disturbances, which is also highly related to the pilot skills. Our soft fingers offer an added layer of production when the fingers collide with the bottom or other obstacles underwater, providing impact absorption for the underwater ROV while providing capable grasping and tactile sensing capabilities. If the original rigid fingers were installed, sudden impacts would occur when a collision happens, causing damage to the robot, the finger and gripper, and the underwater environment.
@@ 5 Discussion
5.1 Encoding Large-Scale, Whole-Body Deformation by Tracking a Single Visual Representation
This study presents a model-based representation by tracking a single visual feature to achieve high-performing reconstruction of large-scale, whole-body deformation for proprioceptive state estimation. We introduced rigidity-aware Aggregated Multi-Handle constraints during the modeling process. This problem is usually characterized by infinite degrees of freedom (DOFs) via a single visual feature in a 6D pose. As a result, we effectively reduced the dimensionality in representing soft, large-scale, whole-body deformation. Our method shows 40 to 700 times faster run-time than commercial software such as Abaqus at different resolutions while exhibiting superior accuracy in deformation reconstruction. Our method also shows 1 to 2 times faster than the widely adopted As-Rigid-As-Possible (ARAP) algorithm. It should be noted that it remains theoretically unsolved to provide a model-based explicit proof regarding this problem, requiring further research in future works. However, our study shows promising capabilities of this approach toward a high-performing solution with real-time reconstruction efficiency and accuracy that can be used for tactile robotics.
5.2 Rigid-Soft Interactive Representation in Tactile Robotics
The guiding principle behind our solution is a physical representation process shared by many existing solutions in Vision-Based Tactile Sensing (VBTS) technologies. Robotics usually interprets the physical world as an object-centric environment, which can be modeled as rigid bodies, soft bodies, or realistic bodies depending on predefined assumptions. A critical task in robotics is to provide a structured, digitalized representation of the unstructured, physical interactions so that the robotic system can make reliable action plans. The various designs of the soft medium in VBTS generally function as a physical filter to transform unstructured, object-centric properties from the external environment into a constrained problem space within the finger towards a refined representation. In this study, we propose a rigid-soft interactive representation using a rigid body (the marker plate) attached to the soft body (the adaptive finger) during contact-based interactions (filled with realistic bodies with various material stiffness). This process is similar to the mass-point model in physics, which provides a succinate placeholder for deriving various physical properties without losing generality in the mathematical formulation. Further development following such representation principle may give researchers a novel perspective to model robotic dynamics as a tactile network of rigid-soft interactive representations, as demonstrated by results reported in this study.
5.3 Vision-Based Multi-Modal Tactile Sensing for Robotics
In this study, we focus our investigation in VBTS on deformation reconstruction only, which can further implement tactile sensing of other perceptual modalities, as demonstrated in our previous work. For example, our recent work [14] achieved state-of-the-art performance in 6D force-and-torque (FT) estimation using a similar design, where a fiducial marker is also attached inside the finger to provide a convenient representation. Combining both methods will achieve a Vision-Based Multi-Modal Tactile Sensing system in our soft robotic finger design, simultaneously providing high-performing tactile sensing in 6D FT and continuous whole-body deformation reconstruction. This will address a significant challenge in robot learning from demonstration [76, 77, 78]. Recent research [79] also shows the possibility of achieving object detection in the external environment using the in-finger vision with a markerless design by implementing the in-painting technique. Our research provides a comprehensive demonstration regarding the robotic potentials of VBTS technology in fundamental theory and engineering applications, contributing to tactile robotics as a promising direction for future research [26].
5.4 Vision-based Tactile Sensing for Amphibious Robotics
Another novelty of this study is the application of VBTS in amphibious robotics. Our study presents comprehensive results and demonstrations in benchmarking performances, shape reconstruction tasks, and system integration with an underwater remotely operated vehicle. Many VBTS solutions require a closed chamber for the miniature camera to implement the photometric principle for tactile sensing, which may become challenging or even unrealistic for a direct application underwater. It should be noted that even after filling the closed chamber with a highly transparent resin to seal the camera, the layer of soft material used on the contact surface needs a depth-dependent calibration that is unrealistic to perform underwater. Furthermore, the soft material, such as silicon gel, will become brittle as the water depth increases [80]. Our previous work already showcased the engineering benefits of our soft robotic finger design, which can be used to reliably estimate 6D FT from on-land to underwater scenarios [74]. In this work, we further demonstrate the applications of VBTS in high-performing shape reconstruction through our soft robotic finger design for amphibious applications. Our soft finger’s metamaterial network leverages structural adaptation by design instead of being solely dependent on the material softness. This significantly reduces the fluidic pressure on our finger’s adaptive behavior. Further discussion of this topic is outside the scope of this study, which we will address in an upcoming work with more details.
6 Conclusion, Limitations, and Future Work
In conclusion, this study presents a novel Vision-Based Tactile Sensing approach for Proprioceptive State Estimation with a focus on amphibious applications. Utilizing a Soft Polyhedral Network structure coupled with marker-based in-finger vision, our method achieves real-time, high-fidelity tactile sensing that accommodates omni-directional adaptations. The introduction of a model-based approach with rigidity-aware Aggregated Multi-Handle constraints enables effective optimization of the soft robotic finger’s deformation. Furthermore, restructuring our proposed approach as an implicit surface model demonstrates superior shape reconstruction and touch-point estimation performance compared to existing solutions. Experimental validations affirm its efficacy in large-scale reconstruction, turbidity benchmarking, and tactile grasping on an underwater Remotely Operated Vehicle, thereby highlighting the potential of tactile robotics for advanced amphibious applications.